OpenStudy (anonymous):

A bus is traveling with a uniform acceleration of 2.75 meters/second2. If the initial velocity of the bus is 16.5 meters/second, what will its velocity be after 9.00 seconds?

OpenStudy (jt950):

u=16.5m/s a=2.75m/s^2 t=9s v=u+at =41.25m/s